# Simulating real-time molecular electron dynamics efficiently using the time-dependent density matrix renormalization group

### v2

@@ Abstract
Compared to ground state electronic structure optimizations, accurate simulations of molecular real-time electron dynamics are usually much more difficult to perform. To simulate electron dynamics, the time-dependent density matrix renormalization group (TDDMRG) has been shown to offer an attractive compromise between accuracy and cost. However, many simulation parameters significantly affect the quality and efficiency of a TDDMRG simulation. So far, it is unclear whether common wisdom from ground state DMRG carries over to the TDDMRG, and a guideline on how to choose these parameters is missing. Here, in order to establish such a guideline, we investigate the convergence behavior of the main TDDMRG simulation parameters, such as time integrator, the choice of orbitals, and the choice of MPS representation for complex-valued non-singlet states. In addition, we propose a method to select orbitals that are tailored to optimize the dynamics. Lastly, we showcase the TDDMRG by applying it to charge migration ionization dynamics in furfural, where we reveal a rapid conversion from an ionized state with a $\sigma$ character to one with a $\pi$ character within less than a femtosecond.
@@ I Introduction
The study of electron dynamics in molecules has become an active research field since the past three decades. [1, 2, 3, 4, 5, 6, 7] This trend has been driven by the development of intense ultrashort pulsed laser systems with wavelengths ranging from as short as those in the X-ray regime up to the infrared.[8, 9, 10] Among others, electron dynamics enables control of chemical reactions or even real-time control of the properties of matter,[11, 12, 13, 14] and a deeper understanding of biomolecules. [3, 15, 5, 16, 17] Alongside the fast progress of experimental methods for ultrafast electron dynamics, accurate theoretical methods that help to interpret experimental observations would be much desirable. In fact, to tackle ultrafast real-time molecular electron dynamics, progress has been made to extend and apply concepts from traditional quantum chemistry. This results in a number of time-dependent extensions of several quantum mechanical methods such as density functional theory (DFT), [18, 19, 20, 21, 22, 23] algebraic diagrammatic construction (ADC), [24, 25, 26] (multireference) configuration interaction (CI), [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37] multiconfiguration self-consistent field (MCSCF)[38, 39, 40, 41, 42] (also known as multi-configuration time-dependent Hartree-Fock, MCTDHF[43, 44, 45, 46, 47, 48, 49, 50, 51]) coupled cluster, [52, 53, 54, 55] density matrix embedding theory,[56] and the density matrix renormalization group (DMRG), [57, 58, 59, 60, 61, 62, 63, 64, 65, 66] to name a few.
The prevailing challenge common to many of these methods is the well-known exponential growth of the computational load as the system size increases. Among these methods, Kohn-Sham real-time time-dependent DFT (RT-TDDFT) [20, 22, 23] can be considered to be the most attractive in terms of computational speed, owing to its underlying assumption in the form of a single electronic configuration. Numerous works employed RT-TDDFT to study the ultrafast electronic motion in molecules would have been too prohibitive for wave function-based methods. RT-TDDFT studies have been able to uncover interesting observations. [67, 68, 69, 70, 6, 71, 72, 73, 74, 75] The main drawbacks of RT-TDDFT, however, are its approximate nature and the difficulty to systematically improve it. This makes it difficult to predict the accuracy of RT-TDDFT results, in particular when the dynamics drives the system far from the ground state. [76, 77, 78, 79, 7, 80]
@@
Time-dependent extensions of the DMRG [83, 84, 85, 86, 87, 88, 89, 90, 91] (TDDMRG [^1] Note that different algorithms exist that exploit DMRG-like approaches for time evolution. Here, we use “TDDMRG” to generically refer to all of them.)[57, 58, 59, 60, 61, 62, 63, 64, 65, 66] offer an attractive compromise—they are based on first principles and allow for increasing the number of orbitals up to a region that would be off-limit for many multireference wave function-based methods. A few studies that demonstrate the capabilities of the DMRG to simulate molecular electron dynamics have been published; among others, these include ionization dynamics, linear response properties, [93] and charge migration.[64, 65] These works included up to $36$ orbitals,
@@
Here, we make an attempt towards a better understanding of many of the required ingredients of time-dependent DMRG simulations for molecular electron dynamics. Our goal is twofold: (1) benchmarking and understanding TDDMRG simulation parameters and (2) extending procedures from ground state DMRG and CAS-based methods for selecting the most suitable orbitals for the active space to dynamics.
Since the interplay of all the algorithms and choices that make up an accurate and efficient DMRG simulation is very complex, this work clearly needs to be limited. In particular, we limit ourselves to only two related TDDMRG variants, namely the time step targeting (TST) method and a projector-splitting approximation of the time-dependent variational principle (TDVP).
Furthermore, we have limited the scope of this paper by only studying the dynamics of one particular physical effects, namely charge migration, which plays a key role in the photofragmentation of peptides [96, 97, 98, 99, 100, 101, 102] and has the potential to enable the control of chemical reactions. [103, 104, 26, 105] Charge migration is initiated through ionization, where a positively charged region, called hole, migrates across the molecule. Notably, this effect happens on an ultrafast scale of attoseconds to few femtoseconds, [106, 3, 4] and is mainly driven by correlation. [1, 107, 108, 2] Thus, its simulation may serve as a measure of how accurate a method is in describing electronic correlation. Several existing time-dependent methods have been used to simulate ultrafast charge migration; these include ADC, [107, 24, 108, 2, 26, 17] RT-TDDFT, [70, 71, 73], time-dependent CI methods, [34, 109, 110, 111] and TDDMRG. [64, 65]
@@ II Theory
Here, we give a concise overview of some necessary ingredients of TDDMRG simulations. For more details, we refer to the references cited in the following sections. In Section II.1 we briefly discuss time-dependent multiconfiguration-based methods, in Section II.2 we review matrix product states, and in Section II.3 the DMRG for computing ground states. We discuss two ways of propagating an MPS in Section II.4 and Section II.5, respectively. This is followed by an overview of how to deal with symmetries in an MPS in Section II.6, and two ways to describe a complex-valued state by an MPS in Section II.7. Lastly, we discuss two ways to make DMRG simulations more efficient, namely the orbital ordering in Section II.8 and general orbital rotations in Section II.9.
@@ II.2 Matrix product states
The matrices appearing in Eq.  (2) are not unique, since one can insert an identity matrix $\mathbf{I}=\mathbf{X}^{-1}\mathbf{X}$ in between any pair of matrices without changing $|{\Psi}\rangle$. This gauge freedom can be exploited to impose an orthonormality condition on $\mathbf{M}^{\sigma_{i}}$. If it satisfies $\sum_{\sigma_{i}}(\mathbf{M}^{\sigma_{i}})^{\dagger}\mathbf{M}^{\sigma_{i}}=\mathbf{I}$, $\mathbf{M}^{\sigma_{i}}\rightarrow\mathbf{L}^{\sigma_{i}}$ and is called left-orthogonalized. If it satisfies $\sum_{\sigma_{i}}\mathbf{M}^{\sigma_{i}}(\mathbf{M}^{\sigma_{i}})^{\dagger}=\mathbf{I}$, $\mathbf{M}^{\sigma_{i}}\rightarrow\mathbf{R}^{\sigma_{i}}$ and is called right-orthogonalized. Using $I-1$ left-orthogonalized sites, matrices $M^{\sigma_{I}}$ at site $I$, and $K-I-1$ right-orthogonalized sites leads to
$|{\Psi}\rangle=\sum_{\{\sigma\}}\left(\prod_{i=1}^{I-1}\mathbf{L}^{\sigma_{i}}\right)M^{\sigma_{I}}\left(\prod_{i=I+1}^{K}\mathbf{R}^{\sigma_{i}}\right)|{\sigma_{1}\ldots\sigma_{K}}\rangle.$ ((3))
Then the norm of the vectorized tensor $M^{[I]}$ is identical to that of the whole MPS. For the form in Eq.  (3), the site $i=I$ is called orthogonality center. Introducing so-called renormalized states $|{l_{\alpha_{I}}}\rangle$ and $|{r_{\alpha_{I}}}\rangle$ of a subsystem formed by the first $I$ sites and the last $K-I-1$ sites, respectively, simplifies Eq.  (3) to
@@ II.4 Time-dependent variational principle
To move the orthogonalization center, the QR decomposition from Eq.  (5) is performed, giving $\mathbf{M}^{[I]}\left(t+\Delta t/2\right)=\mathbf{L}^{[I]}\left(t+\Delta t/2\right)\mathbf{C}^{[I]}\left(t+\Delta t/2\right)$. Before $\mathbf{C}^{[I]}(t+\Delta t/2)$ can be absorbed into $\mathbf{R}^{[I+1]}(t)$ to obtain the new orthogonalization center $\mathbf{M}^{[I]}(t)$, the two propagation times of the tensors need to be aligned. This is achieved by backward-propagating $\mathbf{C}^{[I]}(t+\Delta t/2)$ to time $t$ using Eq.  (7) but with a negative time step. This backward propagation is the only part of the TDVP version of the TDDMRG that does not have a counterpart in the DMRG. Note that the renormalized basis used for backward-propagating $\mathbf{C}^{[I]}(t+\Delta t/2)$ is different from that used for forward-propagating $\mathbf{M}^{[I]}(t)$. This is one of the error sources in TDVP, which can be minimized by decreasing the time step. Performing this sequence of forward and backward propagations for sites $1$ to $K$ and then backward for sites $K$ to $1$ propagates the total MPS from time $t$ to $t+\Delta t$ with an error of $\mathcal{O}(\Delta t^{3})$. As in the DMRG, a two-site version can be introduced as well. Since this algorithm is based on the TDVP, it is often dubbed TDVP-DMRG. It also is known as projector splitting integrator in the multi-layer MCTDH community, where this algorithm is used to propagate tree tensor networks states, an extension of MPSs. [59, 119, 60, 120, 121, 66] For more details, we refer to Refs.60, 61, 63, 66. As with many other propagation methods, eigenstates can be obtained by propagating in imaginary time. Indeed, in the limit of $\Delta t\to-i\infty$, the DMRG algorithm for eigenstates is recovered from the TDVP algorithm.[61]
@@ II.8 Orbital ordering
Due to the linear nature of the MPS, a small perturbation at an orbital/the corresponding site at one end of the MPS needs to “propagate” through the entire MPS before the orbital/site at the other end of the MPS can “respond.” Thus, the MPS is variant under orbital permutations. To achieve optimal performance, the orbitals should be arranged such that those that correlate largely with each other are placed close together. Several orbital ordering methods exist, such as the symmetric reverse Cuthill–McKee ordering, [136] approximate best prefactor ordering using simulated annealing, [137] Fiedler ordering,[138] and using a genetic algorithm to find a close candidate to the global minimum.[139, 88] They are based on minimizing auxiliary quantities such as the bandwidth of the exchange integral matrix. Localized orbitals (see Section II.9) of molecules with simple geometries such as linear ones can also be ordered based on the orbitals’ centers.
@@ III.1 Setup
We choose choloracetylene and furan as test beds. Choloracetylene displays charge migration similar to the well-studied iodoacetylene,[4] and is linear and small enough to allow for the inclusion of all available non-core orbitals of a double-$\zeta$ basis in converged TDDMRG simulations without too large values for the bond dimension. Compared to chloroacetylene, furan is a (slightly) larger prototypical heterocyclic molecule with more complicated charge migration. [110] For furan, we use the 6-31G basis set, [149, 150] resulting in a frozen-core cation full configuration interaction (FCI) space of $48$ and $25$ electrons, while for chloroacetylene we use the def2-SV(P) basis set, [151] resulting in a frozen-core cation FCI space of $41$ orbitals and $15$ electrons. Since our simulations do not target core ionization, we use the frozen core approximation, thus freeze seven (five) orbitals in chloroacetylene (furan). The used geometries are given in Section S1 of the Supporting Information (SI). Note that here, we limit the convergence studies to a particular Gaussian basis size, as our objective is the convergence of TDDMRG and not that of the basis set.
@@
For our choloracetylene simulations, we choose $|{\psi}\rangle$ to be an in-phase $50$/$50$ superposition of the Hartree-Fock HOMO and HOMO-1. Each of these orbitals is a $\pi$ orbital belonging to a $b_{1}$ irrep of the $C_{2v}$ point group. The resulting orbital $|{\psi}\rangle$ is a $p$-like atomic orbital localized at the Cl atom. For our furan simulations, $|\psi\rangle$ is an intrinsic bond orbital (IBO)[154] with a $p$-like symmetry localized at the O atom. All the initial states described above transform according to a particular irrep of the point group of the molecule, and, since the dynamics is field-free, the time evolution conserves the irrep of the initial state. Hence, we make use of the point group symmetry. For chloroacetylene we use the $C_{2v}$ point group, and for furan we use $C_{s}$.

### v3

@@ Abstract
Compared to ground state electronic structure optimizations, accurate simulations of molecular real-time electron dynamics are usually much more difficult to perform. To simulate electron dynamics, the time-dependent density matrix renormalization group (TDDMRG) has been shown to offer an attractive compromise between accuracy and cost. However, many simulation parameters significantly affect the quality and efficiency of a TDDMRG simulation. So far, it is unclear whether common wisdom from ground state DMRG carries over to the TDDMRG, and a guideline on how to choose these parameters is missing. Here, in order to establish such a guideline, we investigate the convergence behavior of the main TDDMRG simulation parameters, such as time integrator, the choice of orbitals, and the choice of matrix product state representation for complex-valued non-singlet states. In addition, we propose a method to select orbitals that are tailored to optimize the dynamics. Lastly, we showcase the TDDMRG by applying it to charge migration ionization dynamics in furfural, where we reveal a rapid conversion from an ionized state with a $\sigma$ character to one with a $\pi$ character within less than a femtosecond.
@@ I Introduction
The study of electron dynamics in molecules has become an active research field for the past three decades. [1, 2, 3, 4, 5, 6, 7] This trend has been driven by the development of intense ultrashort pulsed laser systems with wavelengths ranging from as short as those in the X-ray regime up to the infrared.[8, 9, 10] Among others, electron dynamics enables the control of chemical reactions or even real-time control of the properties of matter,[11, 12, 13, 14] and a deeper understanding of biomolecules. [3, 15, 5, 16, 17] Alongside the fast progress of experimental methods for ultrafast electron dynamics, accurate theoretical methods that help to interpret experimental observations would be much desirable. In fact, to tackle ultrafast real-time molecular electron dynamics, progress has been made to extend and apply concepts from traditional quantum chemistry. This results in a number of time-dependent extensions of several quantum mechanical methods such as density functional theory (DFT), [18, 19, 20, 21, 22, 23] algebraic diagrammatic construction (ADC), [24, 25, 26] (multireference) configuration interaction (CI), [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37] multiconfiguration self-consistent field (MCSCF)[38, 39, 40, 41, 42] (also known as multi-configuration time-dependent Hartree-Fock, MCTDHF[43, 44, 45, 46, 47, 48, 49, 50, 51]) coupled cluster, [52, 53, 54, 55] density matrix embedding theory,[56] and the density matrix renormalization group (DMRG), [57, 58, 59, 60, 61, 62, 63, 64, 65, 66] to name a few.
The prevailing challenge common to many of these methods is the well-known exponential growth of the computational load as the system size increases. Among these methods, the Kohn-Sham real-time time-dependent DFT (RT-TDDFT) [20, 22, 23] can be considered to be the most attractive in terms of computational speed, owing to its underlying assumption in the form of a single electronic configuration. Numerous works employed RT-TDDFT to study the ultrafast electronic motion in molecules that would have been too prohibitive for wave function-based methods. RT-TDDFT studies have been able to uncover interesting observations. [67, 68, 69, 70, 6, 71, 72, 73, 74, 75] The main drawbacks of RT-TDDFT, however, are its approximate nature and the difficulty of systematically improving it. This makes it difficult to predict the accuracy of RT-TDDFT results, in particular when the dynamics drives the system far from the ground state. [76, 77, 78, 79, 7, 80]
@@
Time-dependent extensions of the DMRG [83, 84, 85, 86, 87, 88, 89, 90, 91] (TDDMRG [^1] Note that different algorithms exist that exploit DMRG-like approaches for time evolution. Here, we use “TDDMRG” to generically refer to all of them.)[57, 58, 59, 60, 61, 62, 63, 64, 65, 66] offer an attractive compromise—they are based on first principles and allow for increasing the number of orbitals up to a region that would be off-limit for many multireference wave function-based methods. A few studies that demonstrate the capabilities of the DMRG to simulate molecular electron dynamics have been published; among others, these include ionization dynamics, linear response properties, [93] and charge migration.[64, 65] These works included up to $36$ orbitals.
@@
Here, we make an attempt toward a better understanding of many of the required ingredients of time-dependent DMRG simulations for molecular electron dynamics. Our goal is twofold: (1) benchmarking and understanding TDDMRG simulation parameters and (2) extending procedures from ground state DMRG and CAS-based methods for selecting the most suitable orbitals for the active space to dynamics.
Since the interplay of all of the algorithms and choices that make up an accurate and efficient DMRG simulation is very complex, this work clearly needs to be limited. In particular, we limit ourselves to only two related TDDMRG variants, namely the time step targeting (TST) method and a projector-splitting approximation of the time-dependent variational principle (TDVP).
Furthermore, we have limited the scope of this paper by only studying the dynamics of one particular physical effects, namely charge migration, which plays a key role in the photofragmentation of peptides [96, 97, 98, 99, 100, 101, 102] and has the potential to enable the control of chemical reactions. [103, 104, 26, 105] Charge migration is initiated through ionization, where a positively charged region, called hole, migrates across the molecule. Notably, this effect happens on an ultrafast scale of attoseconds to few femtoseconds, [106, 3, 4] and is mainly driven by correlation. [1, 107, 108, 2] Thus, its simulation may serve as a measure of how accurate a method is in describing electronic correlation. Several existing time-dependent methods have been used to simulate ultrafast charge migration; these include ADC, [107, 24, 108, 2, 26, 17] RT-TDDFT, [70, 71, 73] time-dependent CI methods, [34, 109, 110, 111] and TDDMRG. [64, 65]
@@ II Theory
Here, we give a concise overview of some necessary ingredients of TDDMRG simulations. For more details, we refer to the references cited in the following sections. In Section II.1 we briefly discuss time-dependent multiconfiguration-based methods, in Section II.2 we review matrix product states, and in Section II.3 the DMRG for computing ground states. We discuss two ways of propagating a matrix product state (MPS) in Section II.4 and Section II.5, respectively. This is followed by an overview of how to deal with symmetries in an MPS in Section II.6, and two ways to describe a complex-valued state by an MPS in Section II.7. Lastly, we discuss two ways to make DMRG simulations more efficient, namely the orbital ordering in Section II.8 and general orbital rotations in Section II.9.
@@ II.2 Matrix product states
The matrices appearing in Eq.  (2) are not unique, since one can insert an identity matrix $\mathbf{I}=\mathbf{X}^{-1}\mathbf{X}$ in between any pair of matrices without changing $|{\Psi}\rangle$. This gauge freedom can be exploited to impose an orthonormality condition on $\mathbf{M}^{\sigma_{i}}$. If it satisfies $\sum_{\sigma_{i}}(\mathbf{M}^{\sigma_{i}})^{\dagger}\mathbf{M}^{\sigma_{i}}=\mathbf{I}$, $\mathbf{M}^{\sigma_{i}}\rightarrow\mathbf{L}^{\sigma_{i}}$ and is called left-orthogonalized. If it satisfies $\sum_{\sigma_{i}}\mathbf{M}^{\sigma_{i}}(\mathbf{M}^{\sigma_{i}})^{\dagger}=\mathbf{I}$, $\mathbf{M}^{\sigma_{i}}\rightarrow\mathbf{R}^{\sigma_{i}}$ and is called right-orthogonalized. Using $I-1$ left-orthogonalized sites, matrices $\mathbf{M}^{\sigma_{I}}$ at site $I$, and $K-I-1$ right-orthogonalized sites leads to
$|{\Psi}\rangle=\sum_{\{\sigma\}}\left(\prod_{i=1}^{I-1}\mathbf{L}^{\sigma_{i}}\right)\mathbf{M}^{\sigma_{I}}\left(\prod_{i=I+1}^{K}\mathbf{R}^{\sigma_{i}}\right)|{\sigma_{1}\ldots\sigma_{K}}\rangle.$ ((3))
Then the norm of the vectorized tensor $\mathbf{M}^{[I]}$ is identical to that of the whole MPS. For the form in Eq.  (3), the site $i=I$ is called orthogonality center. Introducing so-called renormalized states $|{l_{\alpha_{I}}}\rangle$ and $|{r_{\alpha_{I}}}\rangle$ of a subsystem formed by the first $I$ sites and the last $K-I-1$ sites, respectively, simplifies Eq.  (3) to
@@ II.4 Time-dependent variational principle
To move the orthogonalization center, the QR decomposition from Eq.  (5) is performed, giving $\mathbf{M}^{[I]}\left(t+\Delta t/2\right)=\mathbf{L}^{[I]}\left(t+\Delta t/2\right)\mathbf{C}^{[I]}\left(t+\Delta t/2\right)$. Before $\mathbf{C}^{[I]}(t+\Delta t/2)$ can be absorbed into $\mathbf{R}^{[I+1]}(t)$ to obtain the new orthogonalization center $\mathbf{M}^{[I]}(t)$, the two propagation times of the tensors need to be aligned. This is achieved by backward-propagating $\mathbf{C}^{[I]}(t+\Delta t/2)$ to time $t$ using Eq.  (7) but with a negative time step. This backward propagation is the only part of the TDVP version of the TDDMRG that does not have a counterpart in the DMRG. Note that the renormalized basis used for backward-propagating $\mathbf{C}^{[I]}(t+\Delta t/2)$ is different from that used for forward-propagating $\mathbf{M}^{[I]}(t)$. This is one of the error sources in TDVP, which can be minimized by decreasing the time step. Performing this sequence of forward and backward propagations for sites $1$ to $K$ and then backward for sites $K$ to $1$ propagates the total MPS from time $t$ to $t+\Delta t$ with an error of $\mathcal{O}(\Delta t^{3})$. As in the DMRG, a two-site version can be introduced as well. Since this algorithm is based on the TDVP, it is often dubbed TDVP-DMRG. It is also known as the projector splitting integrator in the multi-layer MCTDH community, where this algorithm is used to propagate tree tensor networks states, an extension of MPSs. [59, 119, 60, 120, 121, 66] For more details, we refer to Refs.60, 61, 63, 66. As with many other propagation methods, eigenstates can be obtained by propagating in imaginary time. Indeed, in the limit of $\Delta t\to-i\infty$, the DMRG algorithm for eigenstates is recovered from the TDVP algorithm.[61]
@@ II.8 Orbital ordering
Due to the linear nature of the MPS, a small perturbation at an orbital/the corresponding site at one end of the MPS needs to “propagate” through the entire MPS before the orbital/site at the other end of the MPS can “respond.” Thus, the MPS is variant under orbital permutations. To achieve optimal performance, the orbitals should be arranged such that those that correlate largely with each other are placed close together. Several orbital ordering methods exist, such as the symmetric reverse Cuthill–McKee ordering, [136] approximate best prefactor ordering using simulated annealing, [137] Fiedler ordering,[138] and using a genetic algorithm to find a close candidate to the global minimum.[139, 88] They are based on minimizing auxiliary quantities such as the bandwidth of the exchange integral matrix. Localized orbitals (see Section II.9) of molecules with simple geometries such as the linear ones can also be ordered based on the orbitals’ centers.
@@ III.1 Setup
We choose choloracetylene and furan as test beds. Choloracetylene displays charge migration similar to the well-studied iodoacetylene,[4] and is linear and small enough to allow for the inclusion of all available non-core orbitals of a double-$\zeta$ basis in converged TDDMRG simulations without too large values for the bond dimension. Compared to chloroacetylene, furan is a (slightly) larger prototypical heterocyclic molecule with more complicated charge migration. [110] For furan, we use the 6-31G basis set, [149, 150] resulting in a frozen-core cation full configuration interaction (FCI) space of $48$ and $25$ electrons, while for chloroacetylene we use the def2-SV(P) basis set, [151] resulting in a frozen-core cation FCI space of $41$ orbitals and $15$ electrons. Since our simulations do not target core ionization, we use the frozen core approximation, thus freezing seven (five) orbitals in chloroacetylene (furan). The used geometries are given in Section S1 of the Supporting Information (SI). Note that here, we limit the convergence studies to a particular Gaussian basis size, as our objective is the convergence of TDDMRG and not that of the basis set.
@@
For our choloracetylene simulations, we choose $|{\psi}\rangle$ to be an in-phase $50$/$50$ superposition of the Hartree-Fock HOMO and HOMO-1. Each of these orbitals is a $\pi$ orbital belonging to a $b_{1}$ irrep of the $C_{2v}$ point group. The resulting orbital $|{\psi}\rangle$ is a $p$-like atomic orbital localized at the Cl atom. For our furan simulations, $|\psi\rangle$ is an intrinsic bond orbital (IBO)[154] with a $p$-like symmetry localized at the O atom. All of the initial states described above transform according to a particular irrep of the point group of the molecule, and, since the dynamics is field-free, the time evolution conserves the irrep of the initial state. Hence, we make use of the point group symmetry. For chloroacetylene we use the $C_{2v}$ point group, and for furan we use $C_{s}$.